OpenStudy (anonymous):

the area of a right triangle is 40 inches squared angle c is 90 degrees ac is x and bc is x +2 what is x?

OpenStudy (xkehaulanix):

Hmm...not sure if this right or not. But here goes anyways. Area= 40 \[Area =(1/2)(base)(height)\]I'm assuming that ac and bc are referring to the sides, right? And since angle c is the right angle, neither is the longest side of the triangle. So it becomes this: \[40=(1/2)(x)(x+2)\]80=(x)(x+2) \[80=x ^{2}+2x\]\[x^{2} +2x-80=0\](x-8)(x+10)=0 Therefore x= either 8 or -10. Since we are discussing lengths, the negative number cannot be applied. So x is equal to 8.

OpenStudy (anonymous):

Thanks, i must have looked at this thing for an hour and I just wasn't seeing it for some stupid reason

OpenStudy (xkehaulanix):

Haha, no problem. That happens to me a lot too.